# EFX Exists for Three Types of Agents

### v1

@@ Abstract
In this paper, we study the problem of finding an envy-free allocation of indivisible goods among multiple agents. EFX, which stands for envy-freeness up to any good, is a well-studied relaxation of the envy-free allocation problem and has been shown to exist for specific scenarios. For instance, EFX is known to exist when there are only three agents [efx_3], and for any number of agents when there are only two types of valuations [mahara].
In this paper, we show that EFX allocations exist for any number of agents when there are at most three types of additive valuations.
@@ 1 Introduction
Fair division of indivisible resources is a well-researched problem at the intersection of theoretical computer science and economics. The problem arises in a variety of practical settings, from allocating slots or assets to distributing aid or shared goods. One of the most intuitive notions of fairness is envy-freeness (EF) [foley1967resource], where each individual is content with their share compared to others. However, when the resources are indivisible - such as physical objects, housing units, or assets like artwork — achieving true envy-freeness becomes impossible in many cases. While EF provides a natural measure of fairness, the combinatorial nature of indivisible goods often renders EF allocations unattainable, highlighting the necessity for more nuanced fairness criteria.
One such refinement is Envy-Freeness up to any Good (EFX), proposed by Caragiannis et al. [efx_no1], which has gained prominence due to its minimal relaxation of EF. Under EFX, an agent may envy another, but such envy can always be resolved by removing any single good from the envied agent’s bundle. Despite its apparent simplicity, the existence of EFX allocations for an arbitrary number of agents remains an unresolved question. This is true even in many restricted settings, where only partial results are known for specific cases [efx_no1, efx_3, mahara, pr20, ghosal2023efxexistsagentstypes, BergerCFF22].
To discuss EFX allocations in more detail, let us consider the formal framework for discrete fair division. Here, we have a set $\mathcal{N}=\{a_{1},a_{2},\ldots,a_{n}\}$ of $n$ agents and a set =$\mathcal{M}=\{g_{1},g_{2},\ldots,g_{m}\}$ of $m$ indivisible goods. Each agent $a_{i}\in\mathcal{N}$ has a valuation function $v_{i}:2^{\mathcal{N}}\to\mathbb{R}_{\geq 0}$ that quantifies the amount of utility derived by agent $a_{i}$ from any given subset of goods. We focus on additive valuations, which imply that the total value derived by a subset of goods is simply the sum of the values of the individual goods within that subset. Formally, for any subset $S\subseteq\mathcal{M}$, the valuation function can be expressed as: $v_{i}(S)=\sum_{g\in S}v_{i}(\{g\})$. An allocation $X=\langle X_{a_{1}},X_{a_{2}},\ldots,X_{a_{n}}\rangle$ is a partition of the goods into $n$ parts, called as bundles, such that agent $a_{i}$ receives bundle $X_{a_{i}}$.
@@
Given an allocation $X$, an agent $a_{i}$ envies agent $a_{j}$ if $v_{i}(X_{a_{i}})<v_{i}(X_{a_{j}})$. This envy is called weak envy if it disappears after removing any single good from $X_{a_{j}}$. Otherwise, the envy is considered strong. An allocation is said to be envy-free up to any good (EFX) if no agent has strong envy towards any other agent. It is not known whether EFX allocations always exist for $n$ agents with arbitrary additive valuations. However, over the years, several partial results have emerged. Following its introduction by Caragiannis et al. [efx_no1], Plaut and Roughgarden [pr20] showed that EFX allocations exist in two specific cases: $(i)$ when there are only two agents, and $(ii)$ for any number of agents, provided all agents have identical valuations. Mahara [mahara] extended this to show that EFX allocation always exists when each agent is of one of two given types, where agents of the same type have identical additive valuations. In a breakthrough paper, Chaudhury et al. [efx_3] showed that EFX allocations always exist for instances with three agents. Ghosal et al. [ghosal2023efxexistsagentstypes] extended this result to $n$ agents when $n-2$ of them have identical valuations.
@@ 1.1 Our Contributions
The existence of EFX allocations for three types of agents has been an open problem (see, e.g., [Mahara21]). In this work, we resolve this by providing a constructive proof of the following theorem.
@@
EFX allocation always exists for $n$ agents when there are three types of agents, where agents of the same type have identical additive valuations.
@@ 1.2 Related Work
The literature on the fair division of indivisible goods is extensive. This section highlights the works most relevant to our study - EFX allocations of goods; for further details, we refer the reader to a recent survey by Amanatidis et al. [amanatidis2023fair].
In addition to the studies on complete EFX allocations discussed above, there have been several results on almost complete EFX allocations, where the objective is to minimize the number of unallocated goods. A notable contribution by Chaudhury et al. [chaudhury2021little] show that EFX allocations exist with at most $n-1$ unallocated goods, ensuring that no agent envies the set of unallocated goods. Following this, Berger et al. [BergerCFF22] and Mahara [Mahara21] show that EFX allocations can be achieved with at most $n-2$ unallocated goods. Moreover, they show that for four agents, EFX allocations can always be computed with at most one unallocated good.
An interesting constraint is to restrict the number of possible values agents can have for the goods. Amanatidis et al. [amanatidis2021maximum] show that EFX allocations exist for bi-valued instances, where each agent assigns one of two possible values to each good. Garg and Murhekar [garg2023computing] extended this result by showing Pareto optimal EFX allocations always exists for bi-valued instances. In a recent article, Amanatidis et al. [amanatidis2024pushing] show that EFX allocations exist for tri-valued instances.
There have been several more studies on approximate notions of EFX allocations. An allocation is said to be $\alpha$-EFX, if for every pair of agents $a_{i}$ and $a_{j}$, agent $a_{i}$ finds her bundle $X_{a_{i}}$ to be at least as valuable as $\alpha\cdot v_{a_{i}}(X_{a_{j}}\setminus g)$ for any $g\in X_{a_{j}}$. Chan et al. [chan2019maximin] showed the existence of $\frac{1}{2}$-EFX allocations. Following this, Amanatidis et al. [AMANATIDIS202094] showed that $0.618$-EFX allocations always exists. In a recent work, Amanatidis et al. [amanatidis2024pushing] show that $\frac{2}{3}$-EFX allocations exist when there are at most seven agents.
@@ 2 Preliminaries
Let $\mathcal{N}=\{a_{1},a_{2},\dots,a_{n}\}$ be a set of $n$ agents and =$\mathcal{M}=\{g_{1},g_{2},\dots,g_{m}\}$ be a set of $m$ indivisible goods. A discrete fair division instance is represented by the tuple $\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$, where $\mathcal{V}=\langle v_{a_{1}},v_{a_{2}},\dots,v_{a_{n}}\rangle$ comprises the valuation functions. For each agent $a_{i}\in\mathcal{N}$, the function $v_{i}:2^{\mathcal{M}}\to\mathbb{R}_{{\geq}0}$ specifies the utility that agent $a_{i}$ derives from any given subset of goods.
We use the term bundle to denote a subset of goods. An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$ is a tuple of $n$ mutually disjoint bundles such that bundle $X_{a_{i}}$ is assigned to agent $a_{i}$ for all $a_{i}\in\mathcal{N}$. We say an allocation $X$ is complete if $\bigcup_{a_{i}\in\mathcal{N}}X_{a_{i}}=\mathcal{M}$.
Let $a_{i}\in\mathcal{N};\ g\in\mathcal{M};\ S,T\subseteq\mathcal{M}$ and $v_{a_{i}}$ be the valuation function of $a_{i}$. To simplify notation, we write $v_{a_{i}}(g)$ to denote $v_{a_{i}}(\{g\})$ and use $S\setminus g$, $S\cup g$ to denote $S\setminus\{g\}$, $S\cup\{g\}$, respectively. We also write $S>_{a_{i}}T$ to denote $v_{a_{i}}(S)>v_{a_{i}}(T)$ and similarly for $<_{a_{i}},\geq_{a_{i}},\leq_{a_{i}}$ and $=_{a_{i}}$. We use $\min_{a_{i}}(S,T)$ and $\max_{a_{i}}(S,T)$ to denote $\operatorname*{arg\,min}_{Y\in\{S,T\}}v_{a_{i}}(Y)$ and $\operatorname*{arg\,max}_{Y\in\{S,T\}}v_{a_{i}}(Y)$ respectively.
@@
An allocation $Y$ Pareto dominates an allocation $X$, denoted by $Y\succ X$, If for every agent $a_{i}\in\mathcal{N}$ her bundle in $Y$ is at least as good as her bundle in $X$ and for at least one agent, her bundle in $Y$ is strictly better than the one in $X$. Formally, $Y\succ X$ if
$\forall a_{i}\in\mathcal{N},Y_{a_{i}}\geq_{a_{i}}X_{a_{i}}\quad\text{and}\quad\exists a_{j}\in\mathcal{N},Y_{a_{j}}>_{a_{j}}X_{a_{j}}$
Given an allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$, we say that agent $a_{i}$ envies another agent $a_{j}$ if $X_{a_{j}}>_{a_{i}}X_{a_{i}}$. We often also say that agent $a_{i}$ envies the bundle $X_{a_{j}}$. For an allocation $X$, the envy relation among the agents is graphically represented as follows.
@@
Given an allocation $X$, let $E_{X}=(\mathcal{N},E)$ be a directed graph where the set of agents $\mathcal{N}$ is the set of vertices and for every pair $(a_{i},a_{j})$ of agents, $(a_{i},a_{j})\in E$ iff agent $a_{i}$ envies agent $a_{j}$ under the allocation $X$.
The envy from an agent can be further categorized into strong and weak envy as follows.
@@
An envy that is not strong is called as a weak envy. Given the definitions of strong and weak envies, an EFX allocation can be defined as follows.
@@
An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$ is said to be envy free upto any good (EFX) if no agent strongly envies any other agent. That is:
@@
An instance $\mathcal{I}=\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$ is said to be non-degenerate if and only if no agent values two different bundles equally. That is, $\forall{a_{i}\in\mathcal{N}}$ we have $v_{a_{i}}(S)\neq v_{a_{i}}(T)$ for all $S\neq T,\text{ where }S,T\subseteq\mathcal{M}$. Akrami et al. [efx_3_simple] showed that it suffices to deal with non-degenerate instances when there are $n$ agents with general valuation functions, i.e., if each non-degenerate instance has an EFX allocation, each general instance has an EFX allocation.
In the rest of the paper, we only consider non-degenerate instances. This implies that all goods are positively valued by all agents as value of the empty bundle is assumed to be zero. Non-degenerate instances have the following property.
@@
Given any allocation $X$, for any two agents $a$ and $a^{\prime}$ with identical valuation, either $a$ envies $a^{\prime}$ or vice versa.
If an agent $a_{i}$ envies a bundle $S$, we can define a minimally envied subset as follows.
Definition 5 (Minimally envied subset [ChaudhuryKMS21]).
Given an allocation $X$, if an agent $a_{i}$ with bundle $X_{a_{i}}$ envies a bundle $S$, we call $T\subseteq S$ a minimally envied subset of $S$ for agent $a_{i}$ if both the following conditions hold.
  1. 1.
    $X_{a_{i}}<_{a_{i}}T$
  2. 2.
    $X_{a_{i}}\geq_{a_{i}}T\setminus{h}\quad\forall{h\in T}$
For an agent $a_{i}$ and two subsets $S,T\subseteq\mathcal{M}$ such that $S<_{a_{i}}T$, we define $R$ as the minimally envied subset of $T$ with respect to $S$, denoted $R=\mathsf{MES}_{a_{i}}(T,S)$, if $R$ is the subset of $T$ that would be minimally envied by $a_{i}$ assuming $S$ is agent $a_{i}$’s bundle. When $S$ is same as $X_{a_{i}}$, we simply refer to $\mathsf{MES}_{a_{i}}(T,S)$ as $\mathsf{MES}_{a_{i}}(T)$.
Definition 6 (Champion [chaudhury2021little]).
Let $X$ be an allocation and $T\subseteq\mathcal{M}$ be a bundle such that at least one agent envies $T$. An agent $a_{i}$ is said to champion the bundle $T$ if no agent strongly envies the minimally envied subset $\mathsf{MES}_{a_{i}}(T,X_{a_{i}})$.
Definition 7 ($g$-champion [chaudhury2021little]).
@@
Given a partial allocation $X$ along with an unallocated good $g$, a champion graph, denoted by $M_{X}=(\mathcal{N},E_{M})$, is a directed graph with the set of agents $\mathcal{N}$ as the vertices and for every pair of agents $a_{i},a_{j}$, the edge $(a_{i},a_{j})\in E_{M}$ if $a_{i}$ $g$-champions $a_{j}$.
Throughout this paper, “champion edges” refers to $g$-champion edges for some unallocated good $g$, unless otherwise specified. The following is a simple observation about the champion graphs.
@@
For any partial allocation $X$ and corresponding champion graph $M_{X}$ with respect to an unallocated good $g$, $M_{X}$ is always cyclic.
@@
Recall that every agent values every good non-zero. For each agent $a_{i}$, as $a_{i}$ envies $X_{a_{i}}\cup g$ over $X_{a_{i}}$, some agent (possibly $a_{i}$) must $g$-champion $a_{i}$. Thus, every agent has an incoming $g$-champion edge, making $M_{X}$ cyclic. ∎
@@
Given an EFX allocation $X$, which is not necessarily complete, several techniques can be applied to find new EFX allocations that Pareto dominate $X$. We will discuss some of these techniques below.
Firstly, we can eliminate envy cycles.
@@
Chaudhury et al. generalized this idea as follows.
@@
Let $X$ be a partial EFX allocation. If there exists two agents $a_{i}$ and $a_{j}$ such that $a_{i}$ $g$-champions $a_{j}$ and $a_{i}$ is reachable from $a_{j}$ in the envy graph $E_{X}$, then there exists another EFX allocation $Y$ such that $Y\succ X$. Moreover, all the agents in the path from $a_{i}$ to $a_{j}$, including $a_{i}$ and $a_{j}$, get strictly better bundles in $Y$.
@@
Given an allocation $X$, if agent $a_{i}$ does not envy agent $a_{j}$ but $a_{i}$ $g$-champions $a_{j}$, it follows that $g\in\mathsf{MES}_{a_{i}}(X_{a_{j}}\cup g,X_{a_{i}})$. In this case, we say that agent $a_{i}$ decomposes the bundle $X_{a_{j}}$ into top and bottom halves. The top half is defined as $T_{a_{j}}^{a_{i}}=\mathsf{MES}_{a_{i}}(X_{a_{j}}\cup g,X_{a_{i}})\setminus g$, while the bottom half is given by $B_{a_{j}}^{a_{i}}=X_{a_{j}}\setminus T_{a_{j}}^{a_{i}}$. This decomposition allows us to compare the top halves of different agents’ bundles, leading to the following observation.

### v3

@@ Abstract
We study the problem of finding an envy-free allocation of indivisible goods among agents with additive valuations. We focus on the fairness notion of envy-freeness up to any good (EFX). A central open question in fair division is whether EFX allocations always exist for any number of agents. While EFX has been established for three agents [efx_3] and for any number of agents with at most two distinct valuations [mahara], its existence in more general settings remains open.
In this paper, we make significant progress by proving that EFX allocations exist for any number of agents when there are at most three distinct additive valuations. This result simultaneously generalizes both the three-agent case and the two-type case, settling an open question in the field (see [mahara]).
@@ 1 Introduction
Fair allocation of indivisible resources is a fundamental problem at the intersection of economics and computer science. The objective here is to distribute a set of indivisible goods among multiple agents with different preferences in a fair manner. This problem arises in a wide range of practical scenarios, from assigning slots and assets to distributing aid and shared resources. One of the most natural and widely studied fairness notions is envy-freeness (EF) [foley1967resource], which ensures that no agent prefers another’s allocation over their own.
However, envy-free allocations do not always exist, making exact envy-freeness unattainable in many cases. Moreover, even when an EF allocation exists, finding one is computationally intractable. To overcome these challenges, [efx_no1] introduced envy-freeness up to any good (EFX), a minimal and natural relaxation that has gained prominence since then. Under EFX, an agent may envy another, but this envy can always be eliminated by removing any one good from the envied agent’s bundle. EFX is widely regarded as the strongest known fairness criterion that remains potentially achievable for indivisible goods. Quoting [CaragiannisGH19],
Arguably, EFX is the best fairness analog of envy-freeness for indivisible items.
Yet, despite its intuitive appeal, the existence of EFX allocations under general conditions has remained an enigmatic open problem. [procacciaTechnicalperspective20] remarks that:
This fundamental and deceptively accessible question is open. In my view, it is the successor of envy-free cake cutting as fair division’s biggest problem.
To discuss EFX allocations in more detail, we recall the standard framework for discrete fair allocation. The setting consists of a set of $n$ agents, $\mathcal{N}=\{a_{1},a_{2},\dots,a_{n}\}$, and a set of $m$ indivisible goods, =$\mathcal{M}=\{g_{1},g_{2},\dots,g_{m}\}$. Each agent $a_{i}\in\mathcal{N}$ has a valuation function $v_{i}:2^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$, which quantifies the utility that $a_{i}$ derives from any subset of goods. We focus on additive valuations, meaning the value of a subset of goods is simply the sum of the values of its individual items. Formally, for any subset $S\subseteq\mathcal{M}$, the valuation function satisfies $v_{i}(S)=\sum_{g\in S}v_{i}(\{g\})$. An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$ is a partition of $\mathcal{M}$ into $n$ disjoint subsets, called bundles, where each agent $a_{i}$ receives the bundle $X_{a_{i}}$.
@@
Given an allocation $X$, an agent $a_{i}$ envies an agent $a_{j}$ if $v_{i}(X_{a_{i}})<v_{i}(X_{a_{j}})$. This envy is said to be weak if it disappears after removing any one good from $X_{a_{j}}$; otherwise, it is considered strong. An allocation is envy-free up to any good (EFX) if no agent has a strong envy towards any other agent. The question of whether EFX allocations always exist for $n$ agents with arbitrary additive valuations remains open. Following its introduction by [efx_no1], [pr20] proved the existence of EFX allocations in two specific cases: $(i)$ when there are only two agents and $(ii)$ for any number of agents, provided that all agents have identical valuations. [mahara] generalized this result by proving that an EFX allocation always exists when agents belong to one of two types, where agents of the same type have identical additive valuations. In a breakthrough result, [efx_3] established that EFX allocations always exist for instances with three agents. Whether complete EFX allocations exist in significantly more general settings, i.e., either for more number of agents or more than two distinct valuations remains wide open (see, e.g., [mahara]).
@@ 1.1 Our Contributions
In this paper, we significantly advance the state-of-the-art by showing the existence of EFX allocations for instances with three types of agents. Specifically, we provide a constructive proof of the following theorem thereby solving an open question highlighted by [mahara].
@@
An EFX allocation always exists for $n$ agents when there are at most three types of agents, where agents of the same type have identical additive valuations.
Our result generalizes and unifies both the three-agent case [efx_3] and the two-types case [mahara] making a substantial contribution to the field.
Our proof, in addition to using some of the existing techniques in the field, involves several new ideas. We believe that these ideas can be helpful in making further progress on EFX. We outline some of these ideas at a high level in what follows.
Progress Measure:
Our proof is algorithmic in nature and relies on a combination of Pareto improvements and a novel potential function $\phi$, specifically designed for the three-types setting. The potential function is the value of the minimum-valued bundle given to any agent in a fixed group (see Section 3.1 for a formal definition). We start with a trivial partial EFX allocation and iteratively update it. Given any partial EFX allocation, our algorithm constructs a new (possibly partial) EFX allocation with a strictly larger potential. This ensures that the process always terminates with a complete EFX allocation.
Pseudo-Cycles:
All earlier work on EFX allocations have crucially exploited the structural patterns arising from various envy relationships among agents, in order to reallocate goods and improve allocations. Examples of these include envy cycles [lipton] and champion cycles [chaudhury2021little]. In this work, we introduce a novel notion of pseudo-cycles (Section 3.2) that captures certain envy relations among groups of agents. While champion cycles and envy cycles can guide the reallocation of goods and result in Pareto improvement, pseudo-cycles enable us to reallocate goods and increase our potential function, even by possibly assigning lower valued bundles to some agents.
Making Progress Through Regress, and Virtual Bundles:
Our proof goes through an involved case analysis of various configurations of the directed graph of envies among agents (envy graph). In one of the cases, to get a new EFX allocation with a larger value of the potential function $\phi$, we need to iteratively deallocate some goods as an intermediate step. Additionally, we introduce the idea of virtual bundles, where an agent, say $p$, considers a subset of agents and hypothetically assigns smaller virtual bundles to those agents, such that this hypothetical allocation is acceptable to $p$.
For this case, we design an iterative algorithm called the Iterated-Competition algorithm (see Algorithm 1 in Section 6.1). In this algorithm, we use three hierarchical potential functions to track progress. In each iteration of the algorithm, one of the potential functions in the hierarchy improves, while keeping the potential function(s) that are higher up in the hierarchy fixed. The potential functions in the decreasing order of precedence are (1) $\phi$, described earlier, (2) the number of unallocated goods, and (3) the number of virtual bundles. As long as we cannot find an EFX allocation with an improved value of $\phi$, the second potential function is increased while keeping $\phi$ the same. If neither of the two potential functions described above can be increased, then the third potential function increases, keeping the other two fixed.
@@ 1.2 Related Work
The literature on the fair division of indivisible goods is extensive. This section highlights the works most relevant to our study - EFX allocations of goods; for further details, we refer the reader to a recent survey by [amanatidis2023fair].
In addition to the studies on complete EFX allocations discussed above, several works have focused on almost complete EFX allocations, where the objective is to minimize the number of unallocated goods. A key result by [chaudhury2021little] showed that an EFX allocation always exists with at most $n-1$ unallocated goods, such that no agent envies the unallocated set. Subsequent work by [BergerCFF22] and [Mahara21] improved this bound, using much more sophisticated techniques, proving that EFX allocations exist with at most $n-2$ unallocated goods. Additionally, they proved that for four agents, an EFX allocation can always be found with at most one unallocated good. More recently, [ghosal2023efxexistsagentstypes] generalized this result, showing that for any number of agents with at most $k$ distinct valuations, an EFX allocation exists with at most $k-2$ unallocated goods. In particular, the result of [ghosal2023efxexistsagentstypes] implies that one can obtain a partial EFX allocation with at most one unallocated good for the setting with three distinct types of agents. Our result directly improves upon this by obtaining a complete EFX allocation in this setting. In this context, we would like to stress the fact that obtaining complete EFX allocations is significantly more challenging than obtaining partial EFX allocations.
An interesting constraint is to restrict the number of possible values agents can have for the goods. [amanatidis2021maximum] showed that EFX allocations exist for bi-valued instances, where each agent assigns one of two possible values to each good. [garg2023computing] extended this result by showing that Pareto optimal EFX allocations always exist for bi-valued instances.
There have been several more studies on approximate notions of EFX allocations. An allocation is said to be $\alpha$-EFX, if for every pair of agents $a_{i}$ and $a_{j}$, agent $a_{i}$ finds her bundle $X_{a_{i}}$ to be at least as valuable as $\alpha\cdot v_{i}(X_{a_{j}}\setminus g)$ for any $g\in X_{a_{j}}$. [pr20] showed the existence of $\frac{1}{2}$-EFX allocations. Following this, [AMANATIDIS202094] showed that $0.618$-EFX allocations always exists. In a recent work,  [amanatidis2024pushing] showed that $\frac{2}{3}$-EFX allocations exist when there are at most seven agents or when agents have at most three distinct values for the goods.
@@ 2 Preliminaries
Let $\mathcal{N}=\{a_{1},a_{2},\dots,a_{n}\}$ be a set of $n$ agents and =$\mathcal{M}=\{g_{1},g_{2},\dots,g_{m}\}$ be a set of $m$ indivisible goods. A discrete fair division instance is represented by the tuple $\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$, where $\mathcal{V}=\langle v_{1},v_{2},\dots,v_{n}\rangle$ are the valuation functions. For each agent $a_{i}\in\mathcal{N}$, the function $v_{i}:2^{\mathcal{M}}\to\mathbb{R}_{{\geq}0}$ specifies the utility that agent $a_{i}$ derives from any given subset of goods.
Definition 1 (Bundle, Allocation, and Sub-allocation).
We use the term bundle to denote a subset of goods. An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle,$ over the set of agents $\mathcal{N}$ is a tuple of $n$ mutually disjoint bundles, where the bundle $X_{a_{i}}$ is assigned to agent $a_{i}$ for every $a_{i}\in\mathcal{N}$. We say that an allocation $X$ is complete if $\bigcup_{a_{i}\in\mathcal{N}}X_{a_{i}}=\mathcal{M}$; otherwise it is partial.
Given any allocation $X$ and any subset $P\subseteq\mathcal{N}$ of agents, the sub-allocation $X(P)$ is defined as $X(P)=\langle X_{p}\mid p\in P\rangle$, which represents the allocation $X$ restricted to the agents in $P$.
Let $a\in\mathcal{N};\ g\in\mathcal{M};\ S,T\subseteq\mathcal{M}$ and $v_{a}$ be the valuation function of $a$. To simplify notation, we write $v_{a}(g)$ to denote $v_{a}(\{g\})$ and use $S\setminus g$, $S\cup g$ to denote $S\setminus\{g\}$, $S\cup\{g\}$, respectively. We also write $S>_{a}T$ to denote $v_{a}(S)>v_{a}(T)$ and similarly for $<_{a},\geq_{a},\leq_{a}$ and $=_{a}$. We use $\min_{a}(S,T)$ and $\max_{a}(S,T)$ to denote $\operatorname*{arg\,min}_{Y\in\{S,T\}}v_{a}(Y)$ and $\operatorname*{arg\,max}_{Y\in\{S,T\}}v_{a}(Y)$ respectively.
@@
An allocation $Y=\langle Y_{a_{1}},\ldots,Y_{a_{n}}\rangle$ Pareto dominates an allocation $X=\langle X_{a_{1}},\ldots,X_{a_{n}}\rangle$, denoted by $Y\succ X$, if for every agent $a_{i}\in\mathcal{N}$ her bundle $Y_{a_{i}}$ in $Y$ is at least as good as her bundle $X_{a_{i}}$ in $X$, and for at least one agent $a_{j}$, the bundle $Y_{a_{j}}$ is strictly better than $X_{a_{j}}$. Formally, $Y\succ X$ if
$\forall a_{i}\in\mathcal{N},Y_{a_{i}}\geq_{a_{i}}X_{a_{i}}\quad\text{and}\quad\exists a_{j}\in\mathcal{N},Y_{a_{j}}>_{a_{j}}X_{a_{j}}.$
Similarly, given two sub-allocations $Y(P)$ and $X(P)$ over a subset $P\subseteq\mathcal{N}$ of agents, we say $Y(P)\succ X(P)$ if $\forall a_{i}\in P,Y_{a_{i}}\geq_{a_{i}}X_{a_{i}}\quad\text{and}\quad\exists a_{j}\in P,Y_{a_{j}}>_{a_{j}}X_{a_{j}}.$
Given an allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$, we say that an agent $a_{i}$ envies another agent $a_{j}$ if $X_{a_{j}}>_{a_{i}}X_{a_{i}}$. We often also say that the agent $a_{i}$ envies the bundle $X_{a_{j}}$. For an allocation $X$, the envy relation among the agents is represented graphically as follows.
@@
Given an allocation $X$, let $E_{X}=(\mathcal{N},E)$ be a directed graph where the set of vertices of $E_{X}$ is the set of agents $\mathcal{N}$, and for every pair $(a_{i},a_{j})$ of agents, $(a_{i},a_{j})\in E$ iff agent $a_{i}$ envies agent $a_{j}$ under the allocation $X$.
The envy of an agent could either be strong or weak, as defined below:
@@
An envy that is not strong is called a weak envy. Given the definitions of strong and weak envies, an EFX allocation can be defined as follows.
@@
An allocation $X=\langle X_{a_{1}},X_{a_{2}},\dots,X_{a_{n}}\rangle$ is said to be envy-free up to any good (EFX) if no agent strongly envies any other agent. That is,
@@
An instance $\mathcal{I}=\langle\mathcal{N},\mathcal{M},\mathcal{V}\rangle$ is said to be non-degenerate if no agent values two different bundles equally. That is, $\forall{a_{i}\in\mathcal{N}},\ \forall S,T\subseteq\mathcal{M}$, we have $S\neq T\Rightarrow v_{i}(S)\neq v_{i}(T)$. [efx_3_simple] showed that it suffices to deal with non-degenerate instances when there are $n$ agents with general valuation functions, i.e., if each non-degenerate instance has an EFX allocation, each general instance has an EFX allocation. In the rest of the paper, we only consider non-degenerate instances. This implies that all goods are positively valued by all agents as the value of the empty bundle is assumed to be zero. Non-degenerate instances have the following property.
@@
Given any allocation $X$, for any two agents $a$ and $a^{\prime}$ with identical valuation, if at least one of them is allocated a non-empty bundle, then either $a$ envies $a^{\prime}$ or $a^{\prime}$ envies $a$.
For an envied subset of items, we define a minimally envied subset as follows:
Definition 6 (Minimally envied subset [efx_3]).
Given an allocation $X$ and some $S\subseteq\mathcal{M}$ that an agent $a_{i}$ envies, a minimally envied subset of $S$ for agent $a_{i}$ is defined to be a minimum cardinality subset $T$ of $S$ that $a_{i}$ envies. Thus, $(i)\ X_{a_{i}}<_{a_{i}}T$ and $(ii)\ X_{a_{i}}\geq_{a_{i}}T\setminus{h}\quad\forall{h\in T}$.
We also define a minimally envied subset of $S$ for an agent $a_{i}$ with respect to an arbitrary subset of items $R$ analogously. Thus, for an agent $a_{i}$ and $R,S\subseteq\mathcal{M}$ such that $R<_{a_{i}}S$, the set $T$ is said to be a minimally envied subset of $S$ with respect to $R$, denoted $T=\mathsf{MES}_{a_{i}}(S,R)$, if $T$ is the subset of $S$ that would be minimally envied by $a_{i}$ if $R$ is allocated to $a_{i}$. When $R$ is the same as $X_{a_{i}}$, we simply refer to $\mathsf{MES}_{a_{i}}(S,R)$ as $\mathsf{MES}_{a_{i}}(S)$.
Definition 7 (Champion [efx_3]).
Let $X$ be an allocation and $S\subseteq\mathcal{M}$ be envied by at least one agent. An agent $a_{i}$ is said to champion $S$ if $a_{i}=\operatorname*{arg\,min}_{a\in\mathcal{N}}|\mathsf{MES}_{a}(S)|$ i.e. among all the agents that envy $S$, $a_{i}$ is an agent with a minimum cardinality minimally envied subset of $S$.
Definition 8 ($g$-champion [efx_3]).
@@
Given a partial allocation $X$ along with an unallocated good $g$, a champion graph, denoted by $M_{X}^{g}=(\mathcal{N},E_{M})$, is a directed graph with the vertex set same as the set of agents $\mathcal{N}$, and for every pair of agents $a_{i},a_{j}$, the edge $(a_{i},a_{j})\in E_{M}$ if $a_{i}$ $g$-champions $a_{j}$.
The following is a simple observation about champion graphs.
@@
Let $X$ be a partial allocation and $g$ be an unallocated good. Then, the champion graph $M^{g}_{X}$ is always cyclic.
@@
Recall that every agent values every good non-zero. For each agent $a_{i}$, as $a_{i}$ envies $X_{a_{i}}\cup g$ over $X_{a_{i}}$, some agent (possibly $a_{i}$) must $g$-champion $a_{i}$. Thus, every agent has an incoming $g$-champion edge, making $M^{g}_{X}$ cyclic. ∎
For simplicity, we refer to $M_{X}^{g}$ as $M_{X}$ with the implicit understanding that all champion edges correspond to some unallocated good $g$, unless explicitly stated otherwise.
@@
Given an EFX allocation $X$, which is not necessarily complete, several techniques can be applied to find new EFX allocations that Pareto dominate $X$. We will discuss some of these techniques below. Firstly, we can eliminate envy cycles by cyclic exchanges of bundles. This leads to the following lemma:
@@
[chaudhury2021little] generalized this idea as follows.
@@
Let $X$ be a partial EFX allocation. If there exist two agents $a_{i}$ and $a_{j}$ such that $a_{i}$ $g$-champions $a_{j}$, and $a_{i}$ is reachable from $a_{j}$ in the envy graph $E_{X}$, then there exists another EFX allocation $Y$ such that $Y\succ X$. Moreover, all the agents in the path from $a_{i}$ to $a_{j}$, including $a_{i}$ and $a_{j}$, get strictly better bundles in $Y$.
@@
Given an allocation $X$, if agent $a_{i}$ does not envy agent $a_{j}$ but $a_{i}$ $g$-champions $a_{j}$, it follows that $g\in\mathsf{MES}_{a_{i}}(X_{a_{j}}\cup g,X_{a_{i}})$. In this case, we say that agent $a_{i}$ decomposes the bundle $X_{a_{j}}$ into two parts, referred to as the top and bottom halves. The top half is defined as $T_{a_{j}}^{a_{i}}=\mathsf{MES}_{a_{i}}(X_{a_{j}}\cup g,X_{a_{i}})\setminus g$, while the bottom half is given by $B_{a_{j}}^{a_{i}}=X_{a_{j}}\setminus T_{a_{j}}^{a_{i}}$. This decomposition allows us to compare the top halves of different bundles, leading to the following observation.